# Scalable Multi-Robot Collaboration with Large Language Models: Centralized or Decentralized Systems?

### v1

@@ III-A Main Components of LLM Prompt
  • Robot State & Capability: the capabilities (available actions) of each robot and their current location. This prompt component is synthesized by our pre-defined functions. Note that the available actions include possible collisions; it is the responsibility of the planner to find safe plans.
@@ III-C Step History
Including the full history of the dialogue, environment states, and actions rapidly exhausts the context token budget for the LLM planners, constraining the performance of these frameworks. We therefore compare three approaches in an ablation study of the historical information included in the context: (1) no historical information, (2) only state-action pair history (no dialogue), and (3) the full history. We only include the results for all three approaches in the ablation study; since we found that (2) has the best trade-off between task performance and token efficiency (see Section IV-C), all other experiments are performed with only state-action pair history.
@@ III-D Token Length Constraint
We use gpt-4-0613 and gpt-3.5-turbo-0613 in this work, which have context token limits of 8192 and 4097, respectively. To make sure the total token length (prompt + response) does not surpass these limits, we employ a sliding context window over the step history part of the prompt; the step history will include as many of the most recent steps as permissible without surpassing a total prompt length of 3500 tokens.
@@ IV-A Testing Environments
To compare the four different LLM-based planning frameworks, we design four multi-robot task planning environments inspired by a warehouse setting. In order to evaluate how these frameworks scale to many robots, we instantiate each environment with increasing numbers of robots. For BoxNet1 and BoxNet2, we run trials of 4, 8, 16, and 32 robots. For Warehouse and BoxLift, we run trials 4, 6, 8, and 10 robots. For each number of robots in each environment, we perform 10 trials with varied initial conditions, resulting in 40 total trials per environment.
We track whether each trial resulted in task execution. A task is considered a failure in the following conditions: (1) the dialogue among agents results in a context length beyond the token limit, (2) the agents do not reach consensus before a pre-specified limit of dialogue rounds, (3) the syntactic checking iterates beyond a pre-specified limit, (4) the number of planning iterations exceeds a limit before reaching the goal, and (5) the plan results in a collision. We choose the limits for (1), (2), (3), and (4) such that the failure is very likely a result of endless dialogue or actions. Note that only BoxNet2 and Warehouse can have collision.
@@
BoxLift Figure 4(d) shows the BoxLift environment. In this environment, robots are tasked to lift each box (green) in the fewest time steps. The robots are able to lift different amounts of weight, and the boxes have different sizes and weights. In a single time step, multiple robots can be assigned to lift the same box. The box is lifted if the total capability of the robots lifting is greater than the box’s weight. As an additional challenge, the LLM agents are only able to observe the size of each box, not their weight. This is meant to simulate real situations in which box size roughly correlates with weight. The size and weight of each box is roughly proportional, but we introduce some variability. The LLM agents are provided feedback about whether or not the box was successfully lifted. This environment attempts to test the LLM planner’s ability to efficiently assign heterogeneous robots to collaboritve tasks and also incorporate prior experience when planning.
@@ VI CONCLUSION
Our work considers the scalability of LLM-based multi-robot task planning for long-horizon tasks to systems with many robots with heterogeneous capabilities. We propose several new frameworks for collaborative LLM dialogue and find that hybrid approaches with both central and local LLM planners produce the most successful plans and scale best to large number of agents. Future work can explore more complex tasks with more hierarchical frameworks of robot groups, e.g., each agent for each specialized robot sub-group.

### v2

@@ I INTRODUCTION
To the best of our knowledge, this work is the first to study the scalability of different LLM dialogue frameworks for multi-robot planning. The proposed HMAS-2 framework and state-action history method greatly improve the planning success rates for situations under high robot numbers.
@@ III-A Main Components of LLM Prompt
  • Robot State & Capability: the capabilities (available actions) of each robot and their current locations. This prompt part is synthesized by our pre-defined functions. The available actions include possible collisions; it is the responsibility of the planner to find safe plans.
@@ III-C Step History
Including the full history of the dialogue, environment states, and actions rapidly exhausts the context token budget for the LLM planners, constraining the performances. We therefore compare three approaches in an ablation study of the historical information included in the context: (1) no historical information, (2) only state-action pair history (no dialogue), and (3) the full history. We only include the results for all three approaches in the ablation study; since we found that (2) has the best trade-off between task performance and token efficiency (see Section IV-C), all other experiments are performed with only state-action pair history.
@@ III-D Token Length Constraint
We use gpt-4-0613 and gpt-3.5-turbo-0613 in this work, with context token limits of 8192 and 4097, respectively. To make sure the total token length (prompt + response) does not surpass these limits, we employ a sliding context window over the step history part of the prompt; the step history will include as many of the most recent steps as possible without surpassing a total prompt length of 3500 tokens.
@@ IV-A Testing Environments
To compare the four different LLM-based planning frameworks, we design four multi-robot task planning environments inspired by a warehouse setting. In order to evaluate how these frameworks scale to many robots, for BoxNet1 and BoxNet2, we run trials of 4, 8, 16, and 32 robots. For Warehouse and BoxLift, we run trials 4, 6, 8, and 10 robots. For each number of robots in each environment, we perform 10 trials with varied initial conditions.
We track whether each trial resulted in task execution. A task is considered a failure in the following conditions: (1) the dialogue among agents results in a context length beyond the token limit, (2) the agents do not reach consensus before a pre-specified limit of dialogue rounds, (3) the syntactic checking iterates beyond a pre-specified limit, (4) the number of planning iterations exceeds a limit before reaching the goal, and (5) the plan results in a collision. Note that only BoxNet2 and Warehouse can have collision.
@@
BoxLift Figure 4(d) shows the BoxLift environment. In this environment, robots are tasked to lift each box (green) in the fewest time steps. The robots are able to lift different amounts of weight, and the boxes have different sizes and weights. In a single time step, multiple robots can be assigned to lift the same box. The box is lifted if the total capability of the robots lifting is greater than the box’s weight. As an additional challenge, the LLM agents are only able to observe the size of each box, not their weight. This is meant to simulate real situations in which box size roughly correlates with weight. The size and weight of each box is roughly proportional, but we introduce some variability. The LLM agents are provided feedback about whether or not the box was successfully lifted.
@@ VI CONCLUSION
Our work considers the scalability of LLM-based multi-robot task planning for long-horizon tasks to systems with many robots with heterogeneous capabilities. We propose several new frameworks for collaborative LLM dialogue and find that hybrid approaches with both central and local LLM planners produce the most successful plans and scale best to large number of agents. While the experiments focus on warehouse related tasks, the efficient frameworks can generalize to many other multi-robot collaboration situations.
Future work can explore more complex tasks with more hierarchical frameworks of robot groups, e.g., each agent for each specialized robot sub-group. The rapid arising of multi-modal (particularly visual) models can change the communication paradigm for large multi-agent problems, opening up more research topics on agent framework optimization.
VII ACKNOWLEDGEMENTS
This work was supported by ONR under Award N00014-22-1-2478, Army Research Laboratory under DCIST CRA W911NF-17-2-0181, and MIT-IBM Watson AI Lab. This article solely reflects the conclusions of its authors.